# Towards Human-Level 3D Relative Pose Estimation: Generalizable, Training-Free, with Single Reference

### v1

@@ Abstract
Humans can easily deduce the relative pose of a previously unseen object, without labeling or training, given only a single query-reference image pair. This is arguably achieved by incorporating (i) 3D/2.5D shape perception from a single image, (ii) render-and-compare simulation, and (iii) rich semantic cue awareness to furnish (coarse) reference-query correspondence. Existing methods implement (i) by a 3D CAD model or well-calibrated multiple images and (ii) by training a network on specific objects, which necessitate laborious ground-truth labeling and tedious training, potentially leading to challenges in generalization. Moreover, (iii) was less exploited in the paradigm of (ii), despite that the coarse correspondence from (iii) is able to enhance the compare process by filtering out non-overlapped parts under substantial pose differences/occlusions. Motivated by this, we propose a novel 3D generalizable relative pose estimation method by elaborating (i) with a 2.5D shape from an RGB-D reference, (ii) with an off-the-shelf differentiable renderer, and (iii) with semantic cues from a pretrained model like DINOv2. Specifically, our differentiable renderer takes the 2.5D rotatable mesh textured by the RGB and the semantic maps (obtained by DINOv2 from the RGB input), then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, back-propagating the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects, given only a single RGB-D reference, without labeling or training. Extensive experiments on LineMOD, LM-O, and YCB-V show that our training-free method significantly outperforms the state-of-the-art supervised methods, especially under the rigorous Acc@5/10/15 metrics and the challenging cross-dataset settings. The codes are available at https://github.com/ethanygao/training-free_generalizable_relative_pose.
@@ I Introduction
Recent years have witnessed great progress in 3D object pose estimation templates-pose ; nguyen2023nope ; wang2024object ; 3DAHV ; zhao2024dvmnet ; wang2021nemo ; kaushik2024source , which estimates the 3D rotation of an object depicted in a query RGB image. As a key to facilitating interaction with real-world objects, 3D object pose estimation attracts increasing attention from various areas including computer vision, virtual/augmented reality, robotics, and human-computer interaction perez2016robot ; robots ; AR . To date, the community shows great interest in generalizable 3D object pose estimation templates-pose ; onepose ; onepose++ ; relpose ; 3DAHV ; zhao2024dvmnet owing to its wide applicability, which focuses on the generalization to previously unseen objects, preferably in a zero-shot manner [^1] We discuss the relatively easier instance- or category-level object pose estimation in the Related Work Sect. II-A and II-B, respectively..
@@
On the other hand, recent methods propose to reframe the generalizable object pose estimation task as relative pose estimation between a query and a reference images from an unseen object, which is termed as generalizable relative object pose estimation 3DAHV ; zhao2024dvmnet . By treating the reference pose as canonical, estimating the relative pose between the reference-query pair successfully bypasses the laborious 3D scanning (of the CAD reference) or dense views calibration (of the multiple-image reference). However, existing methods rely on a large amount of well-labeled poses between the query-reference pairs to effectively train a neural network, thereby imposing the challenge of acquiring high-quantity training data 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ . Moreover, the generalizability of some network-based methods may be hindered by the training datasets. We empirically find that after pretrained on an external large-scale dataset such as Objaverse objaverse , the current state-of-the-art methods 3DAHV ; zhao2024dvmnet need to perform in-dataset finetune [^2] The in-dataset finetuning denotes that the finetune set comes from the same dataset with the testing set, while not including the testing objects. before testing on the in-dataset unseen objects, which impedes the cross-dataset generalize-ability.
@@
In this context, we work towards universally applicable zero-shot 3D generalizable relative pose estimation, where (i) the object is agnostic/unseen from a cross-dataset, (ii) only a single RGB-D image is available for reference without a 3D CAD model or multi-view images, and (iii) the ground-truth (relative) pose label is not available. In other words, we aim to establish a novel 3D generalizable (in terms of both objects and datasets) relative pose estimation method given only one reference and one query image, without labeling or training. This is extremely challenging due to the mixture of incomplete shape information and missing reference-query correspondence, which leads to a severely degraded optimization problem.
@@
Our method is inspired by the fact that humans can easily infer the relative pose under the aforementioned rigorous setting, even with large pose differences or severe occlusions. We hypothesize that such intelligence is accomplished through (i) perceiving 3D/2.5D shapes from a single image, (ii) conducting render-and-compare simulations via imagination, and (iii) understanding rich semantic cues of the object. For example, given two viewpoints of an unseen animal, humans are able to infer the 3D/2.5D shape of that animal, then identify the correspondences of the animal eyes, noses, ears, etc, and finally rotate and render the 3D/2.5D model until its projection matches the other view. Note that the semantic cues have the potential to deal with the (self-) occluded missing parts, thus enhancing the comparison process, e.g., an animal tail can be simply ignored in the render-and-compare simulations if it only appears in one image and is (self-) occluded in the other.
@@
Specifically, our method starts with an RGB-D reference and an RGB query, where their semantic maps can be obtained by leveraging an advanced pretrained model DINOv2 dinov2 with the RGB inputs [^3] Note that our method possesses the potential of using only an RGB reference, please see the discussion in Sect. I-A (Applicability) and Sect. VI (Limitations and Future Works) for more details. Moreover, our method works reasonably well even without the DINOv2 semantic maps on the LineMOD dataset, as illustrated in Table V.. We leverage an easy-to-use differentiable renderer nvdiffrast nvidiffrast , which takes the RGB- and semantic-textured 2.5D mesh of the reference as input, then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The pose refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, which flows the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects from an arbitrary dataset without labeling or training.
In summary, we propose a novel 3D generalizable relative pose estimation method, which takes only an RGB-D reference and an RGB query pair, without requiring the ground-truth pose labels or training. We achieve this by formulating a pose refinement framework via an off-the-shelf differentiable renderer under the render-and-compare paradigm. Our method does not involve training a network, which naturally possesses zero-shot generalize-ability in terms of both unseen objects and datasets. We conducted extensive experiments on LineMOD linemod , LM-O lmo and YCB-V posecnn datasets. The results from our training-free method exhibit significant improvement over the state-of-the-art supervised methods, e.g., for Acc@15 metric on the LineMOD dataset linemod and the YCB-V dataset posecnn , our label- and training-free method outperforms the supervised state-of-the-art results by 29.98% and 14.28%, respectively.
@@ I-A Taxonomy and Applicability of Our Method
Applicability. Among Table I, the proposed method shares the closest setting to the human intelligence on relative pose estimation that is able to generalize to unseen objects from an arbitrary dataset, with only an additional one-time-collection depth map for the reference image. We have testified in supplementary material that our method can still deliver good estimations with an imprecise depth map, which implies the potential to fully distill human intelligence by a generalizable depth estimator. We note that training a generalizable depth estimator is beyond the scope of, and may introduce distractions to, our current focus. In addition, our method also incorporates the segmentation maps of both query and reference objects as input, which can be obtained by pretrained segmentation models such as SAM sam , FastSAM zhao2023fast and Grounded SAM ren2024grounded . We chose not to delve into these segmentation techniques extensively either, for the same sake of minimizing potential distractions.
@@ II-C Generalizable 6D Pose Estimation
Multi-view reference-based methods can be further divided into feature matching-based and template matching-based approaches. For the former, multi-view reference-based feature matching methods mainly aim to establish 2D-3D correspondences between the RGB query image and sparse point cloud reconstructed by reference views or 3D-3D correspondences between the RGB-D query and RGB-D reference images. For instance, FS6D fs6d designed a dense prototype matching framework by extracting and matching dense RGBD prototypes with transformers. After the correspondence is established, Umeyama umeyama1991least algorithms are utilized for pose estimation. OnePose/Onepose++ onepose ; onepose++ apply the Structure from Motion (SfM) method to reconstruct a sparse point cloud of the unseen object using all reference viewpoints. They then employ an attention-based network to predict the correspondence between 2D pixels and the reconstructed point clouds to estimate the object pose. For the latter, Multi-view references can be reviewed as templates for retrieval when plenty of views exist, or used to reconstruct the 3D object models for template rendering, similar to the CAD-based methods. As an illustration, Gen6D gen6d selects the closest reference view for the query image, and then refines the pose through the 3D feature volume constructed from both the reference and query images. Notably, Gen6D requires more than 200 reference images for initial pose selection. On the contrary, LatentFusion latentfusion reconstructs a latent 3D representation of an object to present an end-to-end differentiable reconstruction and rendering pipeline, and then estimates the pose through gradients update. Since a 3D object representation can be reconstructed utilizing the multi-view information, FoundationPose wen2023foundationpose proposed a unified framework to support both CAD-based and multi-view supported setups. When no CAD model is available, they leverage multi-view references to build a neural implicit representation, which is then used for render-and-compare.
@@ II-D Generalizable Relative Pose Estimation
Recent methods 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ highlight the importance of formulating object pose estimation as a relative pose estimation problem. Specifically, 3DAHV and zhao2024dvmnet address situations where only a single-view reference image is available. 3DAHV evidence that some state-of-the-art feature matching approaches sarlin2020superglue , loftr , zsp fail to generate reliable correspondence between the reference-query pair, while energy-base methods relpose ; relpose++ struggles to capture 3D information. Instead, 3DAHV 3DAHV introduces a framework of hypothesis and verification for generating and evaluating multiple pose hypotheses. Following 3DAHV, DVMNet zhao2024dvmnet directly lifts the 2D image features to 3D voxel information in a hypothesis-free way, computing the relative pose estimation in an end-to-end fashion by aligning the 3D voxels.
Figure 2: The overview of the proposed method. Given an RGB-D reference and an RGB query, we extract the semantic maps from a pretrained DINOv2 model dinov2 for both reference and query. Then, the reference 2.5D front-surface mesh is reconstructed by the depth input without hallucination, which is subsequently texture-mapped by its RGB and semantic images. By leveraging a differentiable renderer nvidiffrast , we generate the rendered RGB and semantic maps using the textured 2.5D reference mesh under a novel view/pose. Finally, the rendered RGB and semantic maps are compared to their query counterparts, producing losses and back-propagating the gradients through the differentiable renderer to refine the relative pose.
@@ III Method
Following the render-and-compare paradigm, current generalizable pose estimation methods often rely on rotatable 3D CAD models or well-calibrated multi-view images, imposing challenges to acquire the 3D CAD models or expensive pose calibration, especially for previously unseen objects. We instead focus on the generalizable relative pose estimation defined in 3DAHV ; zhao2024dvmnet , which aims to estimate the relative pose between a reference-query pair, using only a single reference with an arbitrary pose as canonical (without calibration). Our method differs from 3DAHV ; zhao2024dvmnet in not requiring labeled relative pose to train an estimation network.
@@ III-A Overview
Formally, by using subscript to denote query or reference, our method starts with an RGB pair $I_{r}$ and $I_{q}$ for both reference and query, as well as a depth map $D_{r}$ for the reference. We proposed to estimate the relative pose between $I_{r}$ and $I_{q}$, assisted by $D_{r}$. To this end, we first infer the semantic maps $S_{r}$ and $S_{q}$ from $I_{r}$ and $I_{q}$ exploiting a pretrained DINOv2 model dinov2 . Then, we construct a 2.5D mesh model $M_{r}$ for the reference object based on $D_{r}$, to formulate an RGB and semantic maps textured 2.5D mesh $\mathcal{M}_{r}=\{M_{r},I_{r},S_{r}\}$. Subsequently, the textured 2.5D reference mesh $\mathcal{M}_{r}$ is rotated with an (arbitrary) initial pose $P$ by a differentiable renderer nvidiffrast to generate novel $I_{r}(P)$ and $S_{r}(P)$. Finally, the generated $I_{r}(P)$ and $S_{r}(P)$ are compared with the query $I_{q}$ and $S_{q}$, producing a refinement loss and consequently back-propagate gradients to $P$ through the differentiable renderer. Our method operates the render-and-compare procedure in a self-supervised and network-free manner, without labeling or training.

### v2

@@ Abstract
Humans can easily deduce the relative pose of a previously unseen object, without labeling or training, given only a single query-reference image pair. This is arguably achieved by incorporating i) 3D/2.5D shape perception from a single image, ii) render-and-compare simulation, and iii) rich semantic cue awareness to furnish (coarse) reference-query correspondence. Motivated by this, we propose a novel 3D generalizable relative pose estimation method by elaborating 3D/2.5D shape perception with a 2.5D shape from an RGB-D reference, fulfilling the render-and-compare paradigm with an off-the-shelf differentiable renderer, and leveraging the semantic cues from a pretrained model like DINOv2. Specifically, our differentiable renderer takes the 2.5D rotatable mesh textured by the RGB and the semantic maps (obtained by DINOv2 from the RGB input), then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, back-propagating the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects, given only a single RGB-D reference, without labeling or training. Extensive experiments on LineMOD, LM-O, and YCB-V show that our training-free method significantly outperforms the state-of-the-art supervised methods, especially under the rigorous Acc@5/10/15 metrics and the challenging cross-dataset settings. The codes are available at https://github.com/ethanygao/training-free_generalizable_relative_pose.
@@ I Introduction
Recent years have witnessed great progress in 3D object pose estimation templates-pose ; nguyen2023nope ; wang2024object ; 3DAHV ; zhao2024dvmnet ; wang2021nemo ; kaushik2024source ; gao2019estimation ; gao2016exploiting ; gao2016symmetric , which estimates the 3D rotation of an object depicted in a query RGB image. As a key to facilitating interaction with real-world objects, 3D object pose estimation attracts increasing attention from various areas including computer vision, virtual/augmented reality, robotics, and human-computer interaction perez2016robot ; robots ; AR . To date, the community shows great interest in generalizable 3D object pose estimation templates-pose ; onepose ; onepose++ ; relpose ; 3DAHV ; zhao2024dvmnet owing to its wide applicability, which focuses on the generalization to previously unseen objects, preferably in a zero-shot manner [^1] We discuss the relatively easier instance- or category-level object pose estimation in the Related Work Sect. II-A and II-B, respectively..
@@
On the other hand, recent methods propose to reframe the generalizable object pose estimation task as relative pose estimation between a query and a reference image from an unseen object, which is termed as generalizable relative object pose estimation 3DAHV ; zhao2024dvmnet . By treating the reference pose as canonical, estimating the relative pose between the reference-query pair successfully bypasses the laborious 3D scanning (of the CAD reference) or dense views calibration (of the multiple-image reference). However, existing methods rely on a large amount of well-labeled poses between the query-reference pairs to effectively train a neural network, thereby imposing the challenge of acquiring high-quantity training data 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ . Moreover, the generalizability of some network-based methods may be impeded by the training data. Our empirical findings suggest that after pretrained on an external large-scale dataset such as Objaverse objaverse , the current state-of-the-art methods 3DAHV ; zhao2024dvmnet require in-dataset finetuning [^2] The in-dataset finetuning denotes that the finetune set comes from the same dataset with the testing set, while not including the testing objects. before testing on unseen objects within the dataset, which might potentially hinder their cross-dataset generalizability.
@@
In this context, we work towards universally applicable zero-shot 3D generalizable relative pose estimation, where i) the object is agnostic/unseen from a cross-dataset, ii) only a single RGB-D image is available for reference without a 3D CAD model or multi-view images, and iii) the ground-truth (relative) pose label is not available. In other words, we aim to establish a novel 3D generalizable (in terms of both objects and datasets) relative pose estimation method given only one reference and one query image, without labeling or training. This is extremely challenging due to the mixture of incomplete shape information and missing reference-query correspondence, which leads to a severely degraded optimization problem.
@@
Our method is inspired by the fact that humans can easily infer the relative pose under the aforementioned rigorous setting, even with large pose differences or severe occlusions. We hypothesize that such intelligence is accomplished through i) perceiving 3D/2.5D shapes from a single image, ii) conducting render-and-compare simulations via imagination, and iii) understanding rich semantic cues of the object. For example, given two viewpoints of an unseen animal, humans are able to infer the 3D/2.5D shape of that animal, then identify the correspondences of the animal eyes, noses, ears, etc, and finally rotate and render the 3D/2.5D model until its projection matches the other view. Note that the semantic cues have the potential to deal with the (self-) occluded missing parts, thus enhancing the comparison process, e.g., an animal tail can be simply ignored in the render-and-compare simulations if it only appears in one image and is (self-) occluded in the other.
@@
Specifically, our method starts with an RGB-D reference and an RGB query, where their semantic maps can be obtained by leveraging an advanced pretrained model DINOv2 dinov2 with the RGB inputs [^3] Note that our method possesses the potential of using only an RGB reference, please see the discussion in Sect. I-A (Applicability) and Sect. VI (Limitations and Future Works) for more details. Moreover, our method works reasonably well even without the DINOv2 semantic maps on the LineMOD dataset, as illustrated in Table V.. We leverage an easy-to-use differentiable renderer nvdiffrast nvdiffrast , which takes the RGB- and semantic-textured 2.5D mesh of the reference as input, then renders new RGB and semantic maps (with back-surface culling) under a novel rotated view. The pose refinement loss comes from comparing the rendered RGB and semantic maps with the query ones, which flows the gradients through the differentiable renderer to refine the 3D relative pose. As a result, our method can be readily applied to unseen objects from an arbitrary dataset without labeling or training, and naturally generalizes to cross-dataset scenarios.
In summary, we propose a novel 3D generalizable relative pose estimation method, which takes only an RGB-D reference and an RGB query pair, without requiring the ground-truth pose labels or training. We achieve this by formulating a pose refinement framework via an off-the-shelf differentiable renderer under the render-and-compare paradigm. Our method does not involve training a network, which naturally possesses zero-shot generalizability in terms of both unseen objects and datasets. We conducted extensive experiments on LineMOD linemod , LM-O lmo and YCB-V posecnn datasets. The results from our training-free method exhibit significant improvement over the state-of-the-art supervised methods, e.g., for Acc@15 metric on the LineMOD dataset linemod and the YCB-V dataset posecnn , our label- and training-free method outperforms the supervised state-of-the-art results by 29.98% and 14.28%, respectively. Our contributions are three-fold:
  • We propose a novel and simple relative pose estimation method that naturally generalizes to unseen objects, without the need for ground truth pose labels or neural network training.
  • Our method is optimized by an off-the-shelf differentiable renderer and thus training-free. This eliminates the training-data dependency in supervised CNN/ViT-based methods, thus ensuring inherently robust generalizability.
  • Our method employs a render-and-compare framework leveraging 2.5D meshes, thereby avoiding the challenging 3D hallucination of unseen objects. Building upon this, semantic features (such as those from DINOv2) can be integrated as texture maps (via PCA dimensionality reduction) into the render-and-compare process.
@@ I-A Taxonomy and Applicability of Our Method
Applicability. Among Table I, the proposed method shares the closest setting to the human intelligence on relative pose estimation that is able to generalize to unseen objects from an arbitrary dataset, with only an additional one-time-collection depth map for the reference image.
Our method arguably possesses better applicability compared to the state-of-the-art methods summarized in Table I, as i) unlike supervised in-dataset state-of-the-art methods, our method does not require ground truth (GT) pose annotations for training, where obtaining a large number of GT poses is arguably more challenging than acquiring a single reference depth map. ii) Our method requires only a one-time reference depth annotation, which can be collected and fixed in advance. Moreover, depth sensors are commonly used in our primary application domain, i.e., robotics. We have testified in supplementary material Table S1 that our method can still deliver good estimations with an imprecise depth map, simulating noisy depth sensors. iii) Our training-free method naturally generalizes to unseen objects because it does not involve training a neural network, and thus is training-data independent. In contrast, supervised state-of-the-art methods typically perform less satisfactorily in cross-dataset scenarios, as they suffer from training-data dependency, leading to generalization issues given different training and evaluation datasets.
To further examine the potential of fully distilling human intelligence, we conducted ablation experiments in the supplementary material by substituting the GT depth with predictions from the state-of-the-art Depth Anything v2 yang2024depth . We note that to preserve the object shape, our method requires the metric depth, meaning the estimated depth $z$ and spatial dimensions $x,y$ should share the same unit of measurement (e.g., both in meters). However, as shown in Table S2-S4 in supplementary material, the off-the-shelf metric depth estimator occasionally fails to generalize across different datasets, likely due to an imprecisely recovered depth scale caused by varying camera parameters and/or diverse objects across different training and evaluation datasets. We note that a generalizable metric depth estimator would alleviate this issue, but training a generalizable metric depth estimator is beyond the scope of, and may introduce distractions to, our current focus.
Finally, our method also incorporates the segmentation maps of both query and reference objects as input, which can be obtained by pretrained segmentation models such as SAM sam , FastSAM zhao2023fast and Grounded SAM ren2024grounded . We chose not to delve into these segmentation techniques extensively either, for the same sake of minimizing potential distractions.
@@ II-C Generalizable 6D Pose Estimation
Multi-view reference-based methods can be further divided into feature matching-based and template matching-based approaches. For the former, multi-view reference-based feature matching methods mainly aim to establish 2D-3D correspondences between the RGB query image and sparse point cloud reconstructed by reference views or 3D-3D correspondences between the RGB-D query and RGB-D reference images. For instance, FS6D fs6d designed a dense prototype matching framework by extracting and matching dense RGBD prototypes with transformers. After the correspondence is established, Umeyama umeyama1991least algorithms are utilized for pose estimation. OnePose/OnePose++ onepose ; onepose++ apply the Structure from Motion (SfM) method to reconstruct a sparse point cloud of the unseen object using all reference views. They then employ an attention-based network to predict the correspondence between 2D pixels and the reconstructed point clouds to estimate the object pose. For the latter, multi-view references can be reviewed as templates for retrieval when plenty of views exist, or used to reconstruct the 3D object models for template rendering, similar to the CAD-based methods. Gen6D gen6d selects the closest reference view for the query image, and then refines the pose through the 3D feature volume constructed from both the reference and query images. Notably, Gen6D requires more than 200 reference images for initial pose selection. On the contrary, LatentFusion latentfusion reconstructs a latent 3D representation of an object to present an end-to-end differentiable reconstruction and rendering pipeline, and then estimates the pose through gradients update. Since a 3D object representation can be reconstructed utilizing the multi-view information, FoundationPose wen2023foundationpose proposed a unified framework to support both CAD-based and multi-view supported setups. When no CAD model is available, they leverage multi-view references to build a neural implicit representation, which is then used for render-and-compare.
@@ II-D Generalizable Relative Pose Estimation
Recent methods 3DAHV ; zhao2024dvmnet ; relpose ; relpose++ highlight the importance of formulating object pose estimation as a relative pose estimation problem. Specifically, 3DAHV ; zhao2024dvmnet address situations where only a single-view reference image is available. 3DAHV evidence that some state-of-the-art feature matching approaches sarlin2020superglue ; loftr ; zsp fail to generate reliable correspondence between the reference-query pair, while energy-based methods relpose ; relpose++ struggles to capture 3D information. Instead, 3DAHV 3DAHV introduces a framework of hypothesis and verification for generating and evaluating multiple pose hypotheses. Based on that, DVMNet zhao2024dvmnet directly lifts the 2D image features to 3D voxel information in a hypothesis-free way, computing the relative pose estimation in an end-to-end fashion by aligning the 3D voxels.
Figure 2: The overview of the proposed method. Given an RGB-D reference and an RGB query, we extract the semantic maps from a pretrained DINOv2 model dinov2 for both reference and query. Then, the reference 2.5D front-surface mesh is reconstructed by the depth input without hallucination, which is subsequently texture-mapped by its RGB and semantic images. By leveraging a differentiable renderer nvdiffrast , we generate the rendered RGB and semantic maps using the textured 2.5D reference mesh under a novel view/pose. Finally, the rendered RGB and semantic maps are compared to their query counterparts, producing losses and back-propagating the gradients through the differentiable renderer to refine the relative pose.
@@ III Method
Following the render-and-compare paradigm, current generalizable pose estimation methods often rely on rotatable 3D CAD models or well-calibrated multi-view images, imposing challenges to acquire the 3D CAD models or expensive pose calibration, especially for previously unseen objects. We instead focus on the generalizable relative pose estimation formulated in the pioneering works 3DAHV ; zhao2024dvmnet , which aims to estimate the relative pose between a reference-query pair, using only a single reference with an arbitrary pose as canonical (without calibration). Our method differs from 3DAHV ; zhao2024dvmnet in not requiring labeled relative pose to train an estimation network.
@@ III-A Overview
Formally, by using subscript to denote query or reference, our method starts with an RGB pair $I_{r}$ and $I_{q}$ for both reference and query, as well as a depth map $D_{r}$ for the reference. We proposed to estimate the relative pose between $I_{r}$ and $I_{q}$, assisted by $D_{r}$. To this end, we first infer the semantic maps $S_{r}$ and $S_{q}$ from $I_{r}$ and $I_{q}$ exploiting a pretrained DINOv2 model dinov2 . Then, we construct a 2.5D mesh model $M_{r}$ for the reference object based on $D_{r}$, to formulate an RGB and semantic maps textured 2.5D mesh $\mathcal{M}_{r}=\{M_{r},I_{r},S_{r}\}$. Subsequently, the textured 2.5D reference mesh $\mathcal{M}_{r}$ is rotated with an (arbitrary) initial pose $P$ by a differentiable renderer nvdiffrast to generate novel $I_{r}(P)$ and $S_{r}(P)$. Finally, the generated $I_{r}(P)$ and $S_{r}(P)$ are compared with the query $I_{q}$ and $S_{q}$, producing a refinement loss and consequently back-propagate gradients to $P$ through the differentiable renderer. Our method operates the render-and-compare procedure in a self-supervised and network-free manner, without labeling or training.